Change 25 oz/mL to kg/gal if 1.0 oz = 28.35 g.


1
Expert's answer
2021-05-09T18:17:07-0400
"25\ \dfrac{oz}{mL}\cdot\dfrac{28.35\ g}{1\ oz}\cdot\dfrac{1\ kg}{1000\ g}\cdot\dfrac{1000\ mL}{1\ L}\cdot\dfrac{1\ L}{0.26417\ gal}=2683\ \dfrac{kg}{gal}."